Write each decimal as a fractio or mixed number in lowest term 18.4
Answer by stanbon(75887)

Write each decimal as a fraction or mixed number in lowest term 18.4
--------------------
18.4 = 184/10 = 92/5 = 18 2/5
=================================
Cheers,
Stan H.
============